OpenStudy (anonymous):

Find the differential of the function y=x3sin(9x).

zepdrix (zepdrix):

\[\Large y=x^3\sin9x\]And we're supposed to find the differential dy?

OpenStudy (anonymous):

dx

zepdrix (zepdrix):

dx? :o hmm ok..\[\Large dy=\color{royalblue}{(x^3)'}\sin9x+x^3\color{royalblue}{(\sin9x)'}\]We'll need to apply the product rule. The blue portions we'll need to take derivatives of. Understand the setup? :o

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

3x^2sin9x+x^3cos9x would be right?

zepdrix (zepdrix):

Woops! When you take the derivative of the sin9x, don't forget your chain rule!! :O We should get an extra factor of 9, yes?\[\Large (\sin9x)' \quad=\quad \cos9x(9x)'\quad=\quad 9\cos9x\]Rest of it looks good though.

OpenStudy (anonymous):

wait i am confused

zepdrix (zepdrix):

chain rule confusing?

OpenStudy (anonymous):

yea

zepdrix (zepdrix):

So like.. here's the simpler explanation. We can go with something more formal if this doesn't make sense. Ummmmmm... Chain rule tells us to do a couple things when we take a derivative. Make a copy of the inner function, and multiply by its derivative. Example:\[\Large \frac{d}{dx}(3x-2)^2 \quad=\quad 2(3x-2)\color{royalblue}{\frac{d}{dx}(3x-2)}\] \[\Large =2(3x-2)\color{royalblue}{(3)}\]

zepdrix (zepdrix):

So in that example ^ we applied the power rule to the outermost function (stuff)^2. Then we multiplied by the derivative of the "stuff"

zepdrix (zepdrix):

If you want the more formal definition, you can think of it as a composition of functions. Chain rule says this:\[\Large \left[f(g(x))\right]' \quad=\quad f'(g(x))\cdot g'(x)\] So in our problem we can think of it like this:\[\Large f(x)=\sin x, \qquad g(x)=9x\] \[\Large f(g(x))=\sin9x\]

zepdrix (zepdrix):

\[\Large f'(g(x))=\cos9x,\qquad g'(x)=9\]

zepdrix (zepdrix):

Chain rule can be a tough one to get a good grasp on, even harder to explain lol :(

OpenStudy (anonymous):

thank you for helping me out

zepdrix (zepdrix):

np c:

OpenStudy (cggurumanjunath):

@zepdrix ' i liked your way of clearing the concepts. .